how do you do the elimination method for this question
The school that Stefan goes to is selling tickets to a choral performance. On the first day of ticket sales
the school sold 3 senior citizen tickets and 1 child ticket for a total of $38. The school took in $52 on
the second day by selling 3 senior citizen tickets and 2 child tickets. Find the price of a senior citizen
ticket and the price of a child ticket.

Answer by ankor@dixie-net.com(22740)

let s = price of the senior cit ticket
let c = price of a child's ticket
:
Write an equation for each statement:
:
On the first day of ticket sales the school sold 3 senior citizen tickets and 1 child ticket for a total of $38.
3s + 1c = 38
;
The school took in $52 on the second day by selling 3 senior citizen tickets and 2 child tickets.
3s + 2c = 52
:
Write the two equation for elimination
3s + 2c = 52
3s + 1c = 38
----------------subtraction eliminates s, find c
0 + 1c = $14 for one child ticket
:
Find s using the equation 3s + 1c = 38, replace c with 14
3s + 14 = 38
3s = 38 - 14
3s = 24
s = 24/3
s = $8 for each senior
:
:
You can check these solutions using the 2n equation




F